OpenStudy (peperwqio):

Evaluate the expression 3(7 + 4)^2 − 14 ÷ 7.

alones (alones):

3(112)−\(\frac{ 14 }{ 7 }\) =(3)(121)−\(\frac{ 14 }{ 7 }\) =363−\(\frac{ 14 }{ 7 }\) =363−2 =361

OpenStudy (latinc):

PEMDAS Please (Parenthesis) Excuse (Exponents) My (Multiplication) Dear (Divid) Aunt (Add) Sally (Subtract)

OpenStudy (peperwqio):

I already knew that I was trying to understand it

OpenStudy (latinc):

What don't you understand?

alones (alones):

Damn my exponents tho pepe i was saying 11^2